OpenStudy (anonymous):

Two similar cylinders have heights 6 inches and 10 inches. What is the ratio of their surface areas?

OpenStudy (anonymous):

1 : 4 3 : 5 9 : 10 9 : 25

OpenStudy (anonymous):

3:5

OpenStudy (anonymous):

How do you know?

OpenStudy (anonymous):

two similar cylinders means they have the same radius (r) of each. let us take the heights are h1 and h2 and the surfaces are A1 and A2 A1 = 2 x (pi) x r x h1 A2 = 2 x (pi) x r x h2 A1/A2 = h1/h2 = 6/10 = 3/5

OpenStudy (anonymous):

Ok, Cool, thank you, and welcome to Openstudy. (If you're new) :D

OpenStudy (anonymous):

sure. i just opened an account here :)

OpenStudy (anonymous):

thats wrong